# A NICER view of the â edge-on pulsar PSR J06143329

### v1

@@ 3.3 Priors
Bassa et al. (2016) discovered a WD companion to PSR J06143329. Its observed temperature from optical photometry resulted in an estimated distance in the range âpc, consistent with radio timing parallax measurements (âpc, Shamohammadi et al. 2024, estimated values and uncertainties reported here and all throughout our work are median values with equal-tailed credible intervals.). For our inference, we used a distance prior following a normal distribution , corresponding to the optical photometry constraint as a range, truncated at .
Tight constraints on mass and inclination were obtained by MeerKAT from radio timing thanks to the detection of a strong Shapiro delay (Miles et al., 2025). These resulted in a mass and orbital inclination of respectively444These values are derived from the radio solution provided in the -years MeerKAT PTA data release. and , used in our inference as Gaussian priors, with the sine of the inclination truncated at one. These are consistent with the measurements from Reardon et al. (2023). As in previous X-PSI analyses, we considered the spin axis of PSR J06143329 to be aligned with its orbital momentum, which is expected to be the case at the end of the recycling process (Yang & Li, 2023). The unconstrained inclination hypothesis was also tested and resulted in similar posterior distributions (see Section 4.2).
The interstellar absorption was modeled with the Tbabs555https://pulsar.sternwarte.uni-erlangen.de/wilms/research/tbabs/ model, depending only on the neutral hydrogen column density parameter . The value from HI4PI Collaboration et al. (2016) in the direction of PSR J06143329 is . This value is possibly overestimated as it is integrated along the whole line of sight. Using the relation between radio dispersion measure (DM) and from He et al. (2013) with yields , but this relation is empirical and has a large scatter associated to it. Since there are uncertainties on those measurements, we adopted a wide uniform prior of .
@@ 3.4 Atmosphere model
Knowledge of the atmosphere of the NS is also crucial for PPM, as the inferred radius and geometry can depend on the atmospheric composition (Salmi et al., 2023). The hydrogen atmosphere of the DA spectral type WD companion, from which the NS likely accreted material to spin-up, heavily favors the hydrogen atmosphere hypothesis. Because of gravitational settling, the lighter elements rise to the top of the atmosphere, making the accreted hydrogen the only visible component of the NS atmosphere. Using a model with a different composition implies the assumption of a total depletion of hydrogen in the NS atmosphere. This could in principle happen due to diffuse hydrogen burning, but the timescale for this process to complete is uncertain (Chang & Bildsten, 2003, 2004). We chose to use a fully ionized NSX hydrogen atmosphere (Ho & Lai, 2001) for our exploratory runs, as done in all our previous analyses. A single inference run with a partially ionized NSX atmosphere table (used previously in Miller et al. 2021) was also carried out for our most favored model (see Section 4.2).
@@ 3.5 Background
A phase-independent background is inferred by fitting a background parameter for every instrumental energy channel. Priors of these parameters can be informed by background measurements, models, or left free to vary from up to the total number of observed counts when a more precise prior value is not available. However, to save on the computational expense from sampling hundreds of supplementary parameters, these parameters are marginalized rather than being sampled directly. The procedure is as follows: for a set of source parameters, X-PSI computes the expected source phase-resolved spectrum and a supplementary contribution to the otherwise Poisson likelihood is evaluated. This supplementary contribution is based on the marginalization of background parameters over the background prior, which is computed in practice within a region of the parameter space around the best fitting background parameters (for a precise derivation of this contribution, see Riley 2019, Appendix B.2). These best fitting background parameters associated with the source parameters are the background parameters used for plotting in Section 4.4. This method allows for much faster computations.
Previous works made use of 3C50 background estimates as priors for the NICER background (Salmi et al., 2022; Choudhury et al., 2024a), but we chose here to leave it uninformed and use phase-averaged XMM data instead. The information from the phase-averaged source spectrum derived by XMM allows X-PSI to infer constraints on the source spectrum in the NICER data, hence resulting in NICER background constraints without relying on 3C50 assumptions (see Section 2.1).
The same background marginalization procedure is applied to the phase-averaged XMM data. However, contrary to NICER, the background prior for XMM is tighter as it is informed by the data. As in previous analyses (Riley et al., 2021; Salmi et al., 2022; Choudhury et al., 2024a; Salmi et al., 2024a; Vinciguerra et al., 2024), the backgrounds of the different EPIC instruments were left to vary uniformly in a range around the measured EPIC background spectra, where is the uncertainty on the number of counts generated by a Poisson process (Figure 2). We adopted throughout all our analyses, whereas previous analysis of PSR J07406620 with XMM data used a value of to account for possible biases in the background which was not extracted from the same exposures but rather from blank-sky exposures (Salmi et al., 2022, 2024a). The XMM background spectra for this analysis were extracted from the same exposures as the source spectra, hence this constraint was tightened. This did not prove problematic as, for the EPIC instruments, the contribution of the background to the flux (âcts/bin) is negligible compared to Poisson uncertainty (Figure 2).
@@ 3.6 Instrument responses
We ran joint NICER and XMM fits for all our inference analyses, hence the cross-calibration uncertainties between the instruments had to be taken into account. Each instrument ARF was multiplied by an energy-independent fitted scaling factor parameter. These parameters are , , , , and were all fit using the same uncorrelated Gaussian prior . This choice relaxes assumptions adopted in previous analyses which used a single factor for all XMM instruments.
An exploratory run with uniform priors instead, , resulted in broader scaling factor posteriors still centered around . Interestingly, it showed a strong positive correlation across XMM instruments and no visible correlation with NICER. However, maximum likelihood cross-calibration scaling factors pointed toward the edge of the prior distribution, which is highly unlikely based on current knowledge of the uncertainty on cross-calibration between instruments (, Marshall et al., 2021), hence resulting in biased best fit values. Hence, we opted to keep the Gaussian priors of the scaling factors for all inference runs.
@@ 4 Inference
In this section, we present inference results for the different hot region geometry models that have been tested, and their respective concordance to the data. The taxonomy of different models is the same as in Riley et al. (2019). Results and posterior distributions for all the models can be found in the Zenodo repository.
@@ 4.1 Exploratory ST-U
Model definition â The observed double peaked bolometric pulse profile hints at the presence of at least two hot regions. A simple ST-Uâ model (Single Temperature - Unshared) was tested. It is composed of two non-overlapping circular hot regions with independent positions, sizes and temperatures.
Geometry and modes â The inference analysis results in two modes with similar configurations of the hot regions, where one spot is located near the equator while the other is on either of the rotational poles. This arises from the edge-on viewing geometry of PSR J06143329, which results in a North/South degeneracy on the location of the polar hotspot. Such an effect was already observed during the analysis of PSR J07406620, which is also seen edge-on (Riley et al., 2021; Salmi et al., 2022, 2024a). Both modes have highly similar () marginal 1D posteriors on non geometric parameters, are equally likely and neither is favored by the data.
@@
Inferred parameters â The main peak in the NICER pulse profile (see Section 2.1) at phase (Figure 1) is caused by the equatorial spot, going in and out of view during rotation. The weak peak is caused by the polar hotspot at , which never fully goes out of view. The phases and temperatures of both hotspots are well constrained for both modes (respectively and ). Spot sizes are only constrained to about . The colatitude of the equatorial hotspot is not tightly constrained (). Finally, the radius is rather small and tightly constrained, with km.
Model performance metrics â Values of log-evidence and maximum log-likelihood are given in Table 1. The 2D residuals of all the tested models can be found in the Zenodo repository. ST-Uâ can reproduce the data as the residuals do not feature any significant deviation from random Poisson noise. However, we tested more complex hotspot geometries to check whether it would bring any improvements.
@@ 4.2 Exploratory ST+PDT
The PDT region has 2 possible configurations that explain the data equally well, each with a North/South degeneracy. While both PDT configurations are made of a large and rather cold region (with angular radius and temperature K) along with another small and hot region ( and K), these regions can be either ceding or superseding components. In both configurations, the small and hot regions arrange themselves on the surface to approximately reproduce a similar surface and position of the hotspot as the polar hot region from the ST-Uâ model (see Section 4.4 where a visual representations of similar configurations is provided).
Compared to the posterior predictive ST-Uâ bolometric pulse, the sharpness of the main peak is slightly increased to better reproduce the data. However, the new PDT is not mainly responsible for this pulse and has settled into a configuration where it partly contributes to the main peak while being fully responsible of the weak peak.
@@
Inferred parameters â The inferred radius of km is higher than for ST-Uâ by 1.2âkm, but is still compatible at (Figure 3). However, the radius values also depend on the mode, where the hot superseding configuration corresponds to a radius lower than the hot ceding configuration by 0.5âkm. Comparing mass-radius posteriors, it seems that the ST+PDTâ distribution corresponds to the high radius tail of ST-U. However, this tail was not classified as a distinct ST-Uâ mode by MultiNest.
Model performance metrics â The ST+PDTâ residuals do not feature any deviation from random Poisson noise and is able to explain well the data. The log-evidence of this model is decisively better than the value of the ST-Uâ model using the Kass and Raftery criterion (Kass & and, 1995, here ). Moreover, the max-likelihood also shows strong improvement. All values are reported in Table 1. This makes ST+PDTâ a more suitable model than ST-Uâ to explain the data.
Supplementary runs â We relaxed the orbital angular momentum and spin alignment hypothesis (see Section 3.3) for one ST+PDTâ run by leaving the inclination angle unrestricted. It resulted in three modes, each replicating distinctive geometric spot configuration among the four modes that were seen in the exploratory ST+PDT, but this time with slightly lower inclinations. All modes have inclination and radius constraints consistent with the reference exploratory ST+PDTâ model, with respective inferred values of , and . This indicates that the assumption of alignment of the spin axis with the orbital momentum is not problematic in this case. This run was more computationally expensive while resulting in similar, although not as tight, radius and geometry constraint. Hence, this configuration was only tested for this model.
To assess the effects of a different atmosphere model, an alternative ST+PDT exploratory run was carried out using the partially ionized NSX hydrogen model (see Section 3.3 for the choice of this atmospheric model). We retrieve the same four geometric modes as for the fully ionized atmosphere case with slightly higher, but compatible, equatorial radius constraints (with difference in the median values km). The log-evidence with the partially ionized atmosphere model is better, but not significantly ( Kass & and, 1995). Hence this atmosphere model gives results consistent with those of the fully ionized model, which we use for the rest of the inferences.
@@ 4.3 Exploratory PDT-U
Model definition â A PDT-Uâ model (Protruding Double Temperature - Unshared) is composed of 2 PDT hotspots. Although the ST+PDTÂ results were convincing, it was tested to investigate the possible metric gains that a second PDT hotspot would bring. Note that this model is a much more computationally intensive one than ST-Uâ and ST+PDT.
@@
Inferred parameters â The inferred PDT-Uâ radius km is compatible with the ST+PDTâ one, with a difference in median values of (, Figure 3). Hence, replacing the ST region with a PDT one is not critical for the radius measurement in this case.
Interestingly, for PDT-U, all individual modes yield a higher and than the combined distribution. This was not the case for ST-Uâ or ST+PDT. This happens when the distributions of the different modes intersect only for lower values of and in some parts of the parameter space. Such a phenomenon was already observed when combining inference results from different bursts in Kini et al. (2024).
The posterior distribution of the angular radii of both superseding hotspots spans rather low values around , with the best-fit parameters reaching even lower values with sizes of across both superseding hotspots in both modes.
Model performance metrics â Although PDT-Uâ has a better log-evidence and maximum likelihood than ST+PDTâ (, Table 1), indicative that PDT-Uâ performs better, such a small difference is not decisive according to the Kass and Raftery criteria (Kass & and, 1995). Vinciguerra et al. (2023) previously found that a log-evidence difference of a few units could not tell models apart: in that work, the log-evidence of a more complex model was units greater than for a simpler model, although the data were generated by the simpler model. It is a possibility that the evidence of PDT-Uâ increases, although not significantly, by accumulating evidence in the region of the parameter space where it reproduces ST+PDT.
@@ 4.4 Headline ST+PDT
Model definition â The Headline run was carried out with the ST+PDTâ model, with improved X-PSI resolution settings and sampling hyper-parameters (see values in Appendix A and B) and still with the multi-modal variant of MultiNest. Because the omnipresence of the North/South degeneracy across previous runs, which gave similar radius constraints, slowed down the computation times and made thorough sampling harder, geometric priors were constrained to remove this degeneracy. This was effectively done by forcing the polar PDT region to lie on the North hemisphere of the NS. A test run with the exploratory ST+PDTâ showed that this procedure speeds up computation by a factor of 2 while resulting in similar geometric and radius constraints, with a difference of radius of âm, the same order of the âm Monte-Carlo standard error, an uncertainty inherent to sampling procedures. Moreover, restricting the prior volume allows for more thorough sampling, making it harder to miss high likelihood regions.
Note that the choice of the North hemisphere, instead of the South hemisphere, for the location of the PDT hot region allows for more conservative uncertainties on the radius. Indeed, the two North PDT configurations of the exploratory ST+PDTâ have individual median radii that are respectively the lower and higher values among all the modes found in this exploratory run, with uncertainties greater or equal to that of their southern counterpart.
@@
NICER Background â The background is similar to the one of exploratory runs. However, the deviation of the maximum likelihood background from the inferred distribution median is 3 at energies below âkeV, and is decreasing with increasing energy (right panel of Figure 5). However, note that we do not fit the individual ST, PDT, or background components individually but their sum and the best fit geometry is consistent with the total spectrum expected distribution, which also has smaller uncertainties than individual ST, PDT or background contributions (Figure 5).
@@
Hence, although it is consistent with the posterior geometric distribution, the best fit geometry should be considered with care. Whenever possible, the full posterior distributions of the geometric parameters should be used instead. The maximum a-posteriori parameter vector of the individual modes could also be considered. These can be retrieved in Appendix B (for the geometric parameters) or from the Zenodo repository (for the full parameter vector).
@@
The inferred distribution of reaches the lower bound of the prior (Figure 12). This value is compatible with the HI4PI Collaboration et al. (2016) value, the latter being also compatible with .
A diagnostic to test the convergence of MultiNest is to see whether increasing sampling accuracy affects the posterior distribution and evidence (Higson et al., 2019). Here, the posterior is similar to that of the LR run with the same restricted priors, although uncertainties are slightly larger. This indicates that sampling has likely converged.

### v2

@@ 3.3 Priors
bassa_cool_2016 discovered a WD companion to PSRÂ J06143329. Its observed temperature from optical photometry resulted in an estimated distance in the range âpc, consistent with radio timing parallax measurements (âpc, shamohammadi_meerkat_2024). For our inference, we used a distance prior following a normal distribution , corresponding to the optical photometry constraint as a range, truncated at .
Tight constraints on mass and inclination were obtained by MeerKAT from radio timing thanks to the detection of a strong Shapiro delay (miles_meerkat_2025). These resulted in a mass and orbital inclination of, respectively,333These values are derived from the radio solution provided in the -years MeerKAT PTA data release. and , used in our inference as Gaussian priors, with the sine of the inclination truncated at one. These are consistent with the measurements from reardon_gravitational-wave_2023. As in previous X-PSI analyses, we considered the spin axis of PSR J06143329 to be aligned with its orbital momentum, which is expected to be the case at the end of the recycling process (Yang_inclination_2023). The unconstrained inclination hypothesis was also tested and resulted in similar posterior distributions (see Section 4.2).
The interstellar absorption was modeled with the Tbabs444https://pulsar.sternwarte.uni-erlangen.de/wilms/research/tbabs/ model, depending only on the neutral hydrogen column density parameter . The value from hi4pi_collaboration_hi4pi_2016 in the direction of PSRÂ J06143329 is . This value is possibly overestimated as it is integrated along the whole line of sight. Using the relation between radio dispersion measure (DM) and from he_correlation_2013 with yields , but this relation is empirical and has a large scatter associated with it. Since there are uncertainties on those measurements, we adopted a wide uniform prior of .
@@ 3.4 Atmosphere model
Knowledge of the atmosphere of the NS is also crucial for PPM, as the inferred radius and geometry can depend on the atmospheric composition (salmi_atmospheric_2023). The hydrogen atmosphere of the DA spectral type WD companion, from which the NS likely accreted material to spin-up, heavily favors the hydrogen atmosphere hypothesis. Because of gravitational settling, the lighter elements rise to the top of the atmosphere, making the accreted hydrogen the only visible component of the NS atmosphere. Using a model with a different composition implies the assumption of a total depletion of hydrogen in the NS atmosphere. This could in principle happen due to diffuse hydrogen burning, but the timescale for this process to complete is uncertain (chang_diffusive_2003; chang_evolution_2004). We chose to use a fully ionized NSX hydrogen atmosphere (ho_atmospheres_2001) for our exploratory runs, as done in all our previous analyses. A single inference run with a partially ionized NSX atmosphere table (used previously in miller_radius_2021) was also carried out for our most favored model (see Section 4.2).
@@ 3.5 Background
A phase-independent background is inferred by fitting a background parameter for every instrumental energy channel. Priors of these parameters can be informed by background measurements, models, or left free to vary from up to the total number of observed counts when a more precise prior value is not available. However, to save on the computational expense from sampling hundreds of supplementary parameters, these parameters are marginalized rather than being sampled directly. The procedure is as follows: for a set of source parameters, X-PSI computes the expected source phase-resolved spectrum and a supplementary contribution to the otherwise Poisson likelihood is evaluated. This supplementary contribution is based on the marginalization of background parameters over the background prior, which is computed in practice within a region of the parameter space around the best fitting background parameters (for a precise derivation of this contribution, see riley_neutron_2019, Appendix B.2). These best fitting background parameters associated with the source parameters are the background parameters used for plotting in Section 4.4. This method allows for much faster computations.
Previous works made use of 3C50 background estimates as priors for the NICER background (salmi_radius_2022; choudhury_nicer_2024), but we chose here to leave it uninformed and use phase-averaged XMM data instead. The information from the phase-averaged source spectrum derived by XMM allows X-PSI to infer constraints on the source spectrum in the NICER data, hence resulting in NICER background constraints without relying on 3C50 assumptions (see Section 2.1).
The same background marginalization procedure is applied to the phase-averaged XMM data. However, contrary to NICER, the background prior for XMM is tighter as it is informed by the data. As in previous analyses (riley_nicer_2021; salmi_radius_2022; choudhury_nicer_2024; salmi_radius_2024; vinciguerra_updated_2024), the backgrounds of the different EPIC instruments were left to vary uniformly in a range around the measured EPIC background spectra, where is the uncertainty on the number of counts generated by a Poisson process (Figure 2). We adopted throughout all our analyses, whereas previous analysis of PSR J07406620 with XMM data used a value of to account for possible biases in the background, which was not extracted from the same exposures but rather from blank-sky exposures (salmi_radius_2022; salmi_radius_2024). The XMM background spectra for this analysis were extracted from the same exposures as the source spectra, hence this constraint was tightened. This did not prove problematic as, for the EPIC instruments, the contribution of the background to the flux (âcts/bin) is negligible compared to Poisson uncertainty (Figure 2).
@@ 3.6 Instrument responses
We ran joint NICER and XMM fits for all our inference analyses, hence the cross-calibration uncertainties between the instruments had to be taken into account. Each instrument ARF was multiplied by an energy-independent fitted scaling factor parameter. These parameters are , , , , and were all fit using the same uncorrelated Gaussian prior . This choice relaxes assumptions adopted in previous analyses, which used a single factor for all XMM instruments.
An exploratory run with uniform priors instead, , resulted in broader scaling factor posteriors still centered around . Interestingly, it showed a strong positive correlation across XMM instruments and no visible correlation with NICER. However, maximum likelihood cross-calibration scaling factors pointed toward the edge of the prior distribution, which is highly unlikely based on current knowledge of the uncertainty on cross-calibration between instruments (, marshall_concordance_2021), hence resulting in biased best fit values. Hence, we opted to keep the Gaussian priors of the scaling factors for all inference runs.
@@ 4 Inference
In this section, we present inference results for the different hot region geometry models that have been tested, and their respective concordance to the data. The taxonomy of different models is the same as in riley_nicer_2019. Results and posterior distributions for all the models can be found in the Zenodo repository (zenodo).
@@ 4.1 Exploratory ST-U
Model definition â The observed double peaked bolometric pulse profile hints at the presence of at least two hot regions. A simple ST-Uâ model (Single Temperature - Unshared) was tested. It is composed of two non-overlapping circular hot regions with independent positions, sizes, and temperatures.
Geometry and modes â The inference analysis results in two modes with similar configurations of the hot regions, where one spot is located near the equator while the other is on either of the rotational poles. This arises from the edge-on viewing geometry of PSRÂ J06143329, which results in a North/South degeneracy on the location of the polar hotspot. Such an effect was already observed during the analysis of PSRÂ J07406620, which is also seen edge-on (riley_nicer_2021; salmi_radius_2022; salmi_radius_2024). Both modes have highly similar () marginal 1D posteriors on non geometric parameters, are equally likely, and neither is favored by the data.
@@
Inferred parameters â The main peak in the NICER pulse profile (see Section 2.1) at phase (Figure 1) is caused by the equatorial spot, going in and out of view during rotation. The weak peak is caused by the polar hotspot at , which never fully goes out of view. The phases and temperatures of both hotspots are well constrained for both modes (respectively, and ). Spot sizes are only constrained to about . The colatitude of the equatorial hotspot is not tightly constrained (). Finally, the radius is rather small and tightly constrained, with km. The full ST-Uâ posterior is available in the Zenodo repository (zenodo).
Model performance metrics â Values of log-evidence and maximum log-likelihood are given in Table 1. The 2D residuals of all the tested models can be found in the Zenodo repository (zenodo). ST-Uâ can reproduce the data as the residuals do not feature any significant deviation from random Poisson noise. However, we tested more complex hotspot geometries to check whether it would bring any improvements.
@@ 4.2 Exploratory ST+PDT
The PDT region has two possible configurations that explain the data equally well, each with a North/South degeneracy. While both PDT configurations are made of a large and rather cold region (with angular radius and temperature K) along with another small and hot region ( and K), these regions can be either ceding or superseding components. In both configurations, the small and hot regions arrange themselves on the surface to approximately reproduce a similar surface and position of the hotspot as the polar hot region from the ST-Uâ model (see Section 4.4 where a visual representation of similar configurations is provided).
Compared to the posterior predictive ST-Uâ bolometric pulse, the sharpness of the main peak is slightly increased to better reproduce the data. However, the new PDT is not mainly responsible for this pulse and has settled into a configuration where it partly contributes to the main peak while being fully responsible for the weak peak.
@@
Inferred parameters â The inferred radius of km is higher than for ST-Uâ by 1.2âkm, but is still compatible at (Figure 3). However, the radius values also depend on the mode, where the hot superseding configuration corresponds to a radius lower than the hot ceding configuration by 0.5âkm. Comparing mass-radius posteriors, it seems that the ST+PDTâ distribution corresponds to the high radius tail of ST-U. However, this tail was not classified as a distinct ST-Uâ mode by MultiNest. The full ST+PDTâ posterior is available in the Zenodo repository (zenodo).
Model performance metrics â The ST+PDTâ residuals do not feature any deviation from random Poisson noise and is able to explain well the data. The log-evidence of this model is decisively better than the value of the ST-Uâ model using the Kass and Raftery criterion (kass_bayes_1995, here ). Moreover, the max-likelihood also shows strong improvement. All values are reported in Table 1. This makes ST+PDTâ a more suitable model than ST-Uâ to explain the data.
Supplementary runs â We relaxed the orbital angular momentum and spin alignment hypothesis (see Section 3.3) for one ST+PDTâ run by leaving the inclination angle unrestricted. It resulted in three modes, each replicating distinctive geometric spot configuration among the four modes that were seen in the exploratory ST+PDT, but this time with slightly lower inclinations. All modes have inclination and radius constraints consistent with the reference exploratory ST+PDTâ model, with respective inferred values of , and . This indicates that the assumption of alignment of the spin axis with the orbital momentum is not problematic in this case. This run was more computationally expensive while resulting in similar, although not as tight, radius and geometry constraints. Hence, this configuration was only tested for this model.
To assess the effects of a different atmosphere model, an alternative ST+PDT exploratory run was carried out using the partially ionized NSX hydrogen model (see Section 3.3 for the choice of this atmospheric model). We retrieve the same four geometric modes as for the fully ionized atmosphere case with slightly higher, but compatible, equatorial radius constraints (with difference in the median values km). The log-evidence with the partially ionized atmosphere model is better, but not significantly ( kass_bayes_1995). Hence this atmosphere model gives results consistent with those of the fully ionized model, which we use for the rest of the inferences.
@@ 4.3 Exploratory PDT-U
Model definition â A PDT-Uâ model (Protruding Double Temperature - Unshared) is composed of two PDT hotspots. Although the ST+PDT results were convincing, it was tested to investigate the possible metric gains that a second PDT hotspot would bring. Note that this model is a much more computationally intensive one than ST-Uâ and ST+PDT.
@@
Inferred parameters â The inferred PDT-Uâ radius km is compatible with the ST+PDTâ one, with a difference in median values of (, Figure 3), which makes them consistent with one another.
Interestingly, for PDT-U, all individual modes yield a higher and than the combined distribution. This was not the case for ST-Uâ or ST+PDT. This happens when the distributions of the different modes intersect only for lower values of and in some parts of the parameter space. Such a phenomenon was already observed when combining inference results from different bursts in kini_constraining_2024.
The posterior distribution of the angular radii of both superseding hotspots spans rather low values around , with the best-fit parameters reaching even lower values with sizes of across both superseding hotspots in both modes. The full PDT-Uâ posterior is available in the Zenodo repository (zenodo).
Model performance metrics â Although PDT-Uâ has a better log-evidence and maximum likelihood than ST+PDTâ (, Table 1), indicative that PDT-Uâ performs better, such a small difference is not decisive according to the Kass and Raftery criteria (kass_bayes_1995). vinciguerra_x-psi_2023 previously found that a log-evidence difference of a few units could not tell models apart: in that work, the log-evidence of a more complex model was units greater than for a simpler model, although the data were generated by the simpler model. It is a possibility that the evidence of PDT-Uâ increases, although not significantly, by accumulating evidence in the region of the parameter space where it reproduces ST+PDT.
@@ 4.4 Headline ST+PDT
Model definition â The Headline run was carried out with the ST+PDTâ model, with improved X-PSI resolution settings and sampling hyper-parameters (see values in Appendix A and B) and still with the multi-modal variant of MultiNest. Because the omnipresence of the North/South degeneracy across previous runs, which gave similar radius constraints, slowed down the computation times and made thorough sampling harder, geometric priors were constrained to remove this degeneracy. This was effectively done by forcing the polar PDT region to lie on the North hemisphere of the NS. A test run with the exploratory ST+PDTâ showed that this procedure speeds up computation by a factor of 2 while resulting in similar geometric and radius constraints, with a difference of radius of âm, the same order of the âm Monte Carlo standard error, an uncertainty inherent to sampling procedures. Moreover, restricting the prior volume allows for more thorough sampling, making it harder to miss high likelihood regions.
The choice of the North hemisphere, instead of the South hemisphere, for the location of the PDT hot region allows for more conservative uncertainties on the radius. This is because the two North PDT configurations of the exploratory ST+PDTâ have individual median radii that are, respectively, the lower and higher values among all the modes found in this exploratory run, with uncertainties greater or equal to that of their southern counterpart.
@@
NICER Background â The background is similar to the one of exploratory runs. However, the deviation of the maximum likelihood background from the inferred distribution median is 3 at energies below âkeV, and is decreasing with increasing energy (right panel of Figure 5). However, note that we do not fit the individual ST, PDT, or background components individually but their sum and the best fit geometry are consistent with the total spectrum expected distribution, which also has smaller uncertainties than individual ST, PDT or background contributions (Figure 5).
@@
Hence, although it is consistent with the posterior geometric distribution, the best fit geometry should be considered with care. Whenever possible, the full posterior distributions of the geometric parameters should be used instead. The maximum a-posteriori parameter vector of the individual modes could also be considered. These can be retrieved in Appendix B (for the geometric parameters) or from the Zenodo repository (for the full parameter vector, zenodo).
@@
The inferred distribution of reaches the lower bound of the prior (see Appendix B for the posterior distribution of a subset of parameters, the full Headline ST+PDTâ posterior is available in the Zenodo repository from zenodo). This value is compatible with the hi4pi_collaboration_hi4pi_2016 value, the latter being also compatible with .
A diagnostic to test the convergence of MultiNest is to see whether increasing sampling accuracy affects the posterior distribution and evidence (higson2019diagnostic). Here, the posterior is similar to that of the LR run with the same restricted priors, although uncertainties are slightly larger. This indicates that sampling has likely converged.